Give an example of an infinite set with finite number of limit points, with proper justification.

1
Expert's answer
2022-01-20T17:15:11-0500

"\text{The set}\\
\Big\{\frac{1}{n} \mid n \in \mathbb{Z}^+\Big\} \bigcup \\ \Big\{1+ \frac{1}{n} \mid n \in \mathbb{Z}^+\Big\} \\ 
\text{ is an infinite set which has a finite limit point} \\
\text{By the definition of limit, we can write}\\
\Big\{\frac{1}{n} \mid n \in \mathbb{Z}^+\Big\} \\ \text{as}\\
\lim_{n \to \infty} \frac{1}{n} = 0 \\
\text{while for}\\
\Big\{1+ \frac{1}{n} \mid n \in \mathbb{Z}^+\Big\}\\
\text{write as}\\
\lim_{n \to \infty} (1 + \frac{1}{n}) = 1 \\
\text{Hence,}\\
\Big\{\frac{1}{n} \mid n \in \mathbb{Z}^+\Big\} \bigcup \\ \Big\{1+ \frac{1}{n} \mid n \in \mathbb{Z}^+\Big\} = \{0,1\} \, \text{which is finite.}"